# Testing Monotonicity of Real-Valued Functions on DAGs

### v1

@@ 1.3.3 Upper bounds
Bottleneck routing from a violation matching (Section 5).
Our algorithmic upper bounds are guided by a single structural fact: distance to monotonicity is witnessed by a large matching of violated comparable pairs in the transitive closure. Concretely, if $f$ is $\varepsilon$-far, then the violation graph (on comparable pairs) contains a matching $M$ of size $|M|\geq\varepsilon n$ (via Konig’s theorem), so there are many disjoint source–sink violations $(u\!\rightsquigarrow\!v,\ f(u)>f(v))$. The difficulty is that these violations typically live in the transitive closure $\mathrm{TC}(G)$, while the input provides only the transitive reduction (Hasse diagram), whose edge set can be much smaller and may contain very few direct violations. The message of Section 5 is that this apparent obstacle can be turned into a win: if direct violating edges are scarce, then they must act as bottlenecks through which many matched violations are routed, and this forces a large supply of violated pairs that can be detected by simple sampling.
The core combinatorial viewpoint is a bottleneck partition of the violation matching $M$. Fix, for each $(u,v)\in M$, an arbitrary directed path $P_{uv}$ in the transitive reduction from $u$ to $v$. Since $f(u)>f(v)$, along that path there must be a first edge $(x,y)$ where the inequality flips, i.e., a violating edge of the reduction. Assign $(u,v)$ to this first violating edge, and thereby partition
@@
This routing/assignment does two crucial things simultaneously.
@@
    Because $M$ is a matching, the sets of left endpoints $\{A_{e}\}$ and right endpoints $\{B_{e}\}$ of the parts $M_{e}$ are disjoint across different bottlenecks $e$. This disjointness is what makes the probabilistic analysis clean and is the main technical reason the approach works at scale.
From here, Section 5 develops two testers (Algorithms 1 and 2), each exploiting a different way to capitalize on this dichotomy and proved in Theorems 5.1 and 5.2.
@@
The first tester (Algorithm 1, see also Section 5.1) samples a moderate number of edges from the transitive reduction $E$ and, independently, a moderate number of comparable pairs from $E(\mathrm{TC}(G))$. The analysis is built around a sharp either–or principle:
  • If many reduction edges are violated (i.e., $|F_{f}|$ is large), then uniform sampling from $E$ finds a witness quickly.
  • If few reduction edges are violated, then the bottleneck partition above implies a quadratic blow-up in the number of violated comparable pairs in the closure: each bottleneck edge $e$ with a large bucket $M_{e}$ certifies many violated pairs of the form $(a,b)$ with $a\in A_{e}$, $b\in B_{e}$. The median trick used in the proof formalizes this amplification: by splitting endpoints around a median threshold $t_{e}$, one obtains large subsets $A^{+}_{e}\subseteq A_{e}$, $B^{-}_{e}\subseteq B_{e}$ such that every $(a,b)\in A^{+}_{e}\times B^{-}_{e}$ is a violated pair in $\mathrm{TC}(G)$. Thus each $e$ contributes $\Omega(|M_{e}|^{2})$ violated comparable pairs, and summing (plus Cauchy–Schwarz) yields a lower bound of order $|M|^{2}/|F_{f}|$ on the number of violated closure edges.
@@
samples in each loop (up to constants). Conceptually, this is a tradeoff between easy local witnesses and many global witnesses: if one cannot see violations locally, they must manifest globally in abundance, and vice versa. This bound can beat the classical $O(\sqrt{n/\varepsilon})$ tester on sparse reachability structures, precisely when $m\ell=o(n^{3})$ for constant $\varepsilon$.
@@
The second tester (Algorithm 2, see also Section 5.2) removes dependence on $\ell$ altogether by avoiding uniform sampling from $\mathrm{TC}(G)$. Instead, after a first stage that samples reduction edges (to catch the many-violating-edges case), it runs a second stage that samples two random multisets of vertices $Q_{L},Q_{R}$ and then checks all comparable pairs between them (reachability tests and scanning $Q_{L}\times Q_{R}$ are free in the model). This is where the bottleneck partition becomes even more powerful: the disjoint endpoint structure means that if there are few violating edges, then the matching $M$ must be split across few buckets, so some bucket $M_{e}$ must be large; and if $Q_{L}$ hits $A^{+}_{e}$ and $Q_{R}$ hits $B^{-}_{e}$ for the same $e$, the tester certifiably rejects.
The technical obstacle in making this quantitative is dependence: sampling vertices with replacement creates complicated correlations when analyzing the event “for some $e$, we hit both sides of $A^{+}_{e}\times B^{-}_{e}$.” Section 5 resolves this with a classic but elegant tool, Poissonization/Poisson approximation. By coupling the multinomial sampling counts with independent Poisson variables, the analysis turns the success events for different bottlenecks $e$ into (essentially) independent events, exactly because the $A^{+}_{e}$’s are disjoint and the $B^{-}_{e}$’s are disjoint. This produces a clean product-form bound and again leverages Cauchy–Schwarz to relate $\sum_{e}|M_{e}|^{2}$ to $|M|^{2}/|F_{f}|$. Choosing the stage sizes to balance the many-violated-reduction-edges regime with the few-bottlenecks regime yields
@@ 1.4 Related work
Monotonicity testing on the Boolean hypercube was initiated by Goldreich et al. [21], with improved testers and near-optimal bounds for hypercubes and hypergrids in subsequent work [11, 12, 16]; connections to isoperimetric inequalities for the Boolean cube were developed in [26]. A substantial body of work also established lower bounds over product domains and hypergrids [5, 10, 13, 14, 15]. In parallel, for general posets, testing was initiated by Fischer et al. [19], and a complementary line uses transitive-closure spanners (TC-spanners) and reachability sparsification to design testers on structured graph families [7, 28]. Beyond standard testing, distance estimation and tolerant testing were studied in [1, 27], while local reconstruction of monotone functions and its limitations appear in [6, 31]. More recently, in product domains, real-valued monotonicity has been connected to isoperimetric inequalities for ordered ranges [9, 12]. Against this backdrop, for explicit-DAG monotonicity testing with real-valued ranges, prior lower bounds in both non-adaptive and adaptive models were far from the $\sqrt{n}$ scale.
@@
Graphs whose edges can be partitioned into induced matchings (Ruzsa–Szemerédi-type graphs) have been extensively studied in extremal graph theory since [30]. In particular, when each induced matching has linear size $r=cn$, the number of matchings in such a decomposition is known to be severely limited (up to constants/logarithmic factors depending on $c$), highlighting a barrier for RS-based lower-bound templates [3, 20]. Positive matchings have also been studied under the name of positive matching decompositions [18]. The graph-theoretic motivation is that positivity gives a structured class of matchings, so one can ask whether the whole edge set can be partitioned into positive matchings and how many parts are necessary. Concretely, positive matchings are characterized by the absence of $M$-alternating closed walks (where $M$ denotes the matching under consideration) on the $M$-saturated vertices [18]; in bipartite graphs this is equivalent to being alternating-cycle-free (uniquely restricted matching) [24]. Our PMRS framework can be viewed as a packing/RS-style relaxation tailored to real-valued monotonicity testing on bipartite DAGs, where positivity is exactly the realizability condition for “violations supported on a matching.”
@@
Section 2 introduces notation and basic facts. Section 3 develops the non-adaptive two-sided lower bounds and explicit PMRS constructions. Section 4 proves adaptive one-sided lower bounds. Section 5 collects bounds parameterized by $m$ and $\ell$, including the algorithmic upper bounds and an $m$-dependent two-sided lower bound.
@@ 3.3 Non-adaptive two-sided lower bounds from PMRS
We begin by proving a non-adaptive two-sided lower bound from PMRS: for any $(s,4\varepsilon)$-PMRS graph, any non-adaptive two-sided tester requires $\Omega(\sqrt{s})$ queries. The proof builds a YES/NO distribution using a margin-normalized witness and a correlation gadget, which we set up below.
@@ 3.5 Main lower bound
We now combine the PMRS lower-bound template with the explicit construction. As a first step, we state a refinement lemma that converts larger positive matchings into many smaller ones while preserving positivity.
@@ 3.6 Bounded ordered ranges and polylogarithmic ranges
The PMRS lower-bound template does not inherently require unboundedness (two-sided).
The non-adaptive two-sided lower bound derived from an $(s,\varepsilon_{0})$-PMRS decomposition (Theorems 3.8 and 3.5) is information-theoretic: it constructs YES/NO distributions by adding a constant-alphabet noise to a slack monotone base and argues that a non-adaptive query set cannot distinguish the two distributions unless it queries both endpoints of some matching edge. The proof uses only the relative order of function values and a constant slack/margin separation; it does not rely on the range being unbounded beyond the ability to realize the prescribed pattern of inequalities. Consequently, the same template extends to $[B]$ provided that the construction can be implemented with values lying in $[B]$ (after an order-preserving relabeling and an affine scaling/translation if needed).
@@ 4 Adaptive One-Sided Lower Bounds
The rest of this section is organized as follows. We first set up the PMRS scaffold and truncated Gibbs hard distribution (Section 4.1), prove edge-tail bounds and parameter admissibility (Section 4.2), establish farness and adaptive spurious-control (Section 4.3), and bound adaptive information leakage via a Green-function comparison (Section 4.4). Next, we combine these ingredients with decoding and Fano’s inequality (Section 4.5) to obtain a fixed-budget barrier (Section 4.6). Finally, we instantiate parameters on an explicit dense scaffold to derive the $\Omega(\sqrt{n})$ adaptive one-sided lower bound (Section 4.7).
@@
(A) Global parameters and scaffold quantities.
@@
$n,n_{0}$ opening of this section total number of vertices and side size in the bipartite scaffold ($n=2n_{0}$) throughout this section
@@
$\delta,\Delta,m$ opening definitions in Section 4.1 min/max degrees and edge count of the scaffold graph Sections 4.2 and 4.7
$Q$ opening definitions in Section 4.1 query-budget scale for leakage and budget barriers Sections 4.4 and 4.6
@@
4.1 PMRS scaffold and truncated soft Gibbs model
In this subsection, we define the truncated Gibbs hard distribution on a PMRS scaffold and record the basic one-site curvature setup used later.
@@ 4.1 PMRS scaffold and truncated soft Gibbs model
Step 5: keep an explicit prefactor. Using $m_{v}=\alpha+2\lambda\deg(v)\geq 2\lambda\deg(v)$ and, since $\deg(v)\geq\delta$ and we assume $\alpha\leq(\lambda+\beta)\delta$, we have $M_{v}\leq\alpha+2(\lambda+\beta)\deg(v)\leq 3(\lambda+\beta)\deg(v)$, together with $A\geq c_{1}\deg(v)(\beta-\lambda)\gamma$, we obtain
@@ 4.2 Edge-tail bounds and admissible parameter set
Standing structural hypothesis.
The next lemma (and only this lemma) uses an additional structural property of the scaffold graph: we assume that the underlying undirected graph of $U$ is $C_{4}$-free (has no $4$-cycles). This hypothesis is satisfied by the explicit family constructed in Section 4.7 below. All subsequent uses of $p_{\mathrm{drift}}(u)$ and the edge-tail bounds depend on this hypothesis.
@@
4.3 Farness and adaptive spurious-control
@@ 4.6 Budgeted one-sided lower bound
We now combine the spurious-control and leakage bounds from the previous subsections to obtain a fixed-budget barrier for adaptive one-sided testers.
Theorem 4.22 (Fixed-budget information barrier).
@@
4.7 A $\Omega(\sqrt{n})$ instantiation and closing mean control
To apply the fixed-budget barrier Theorem 4.22, we verify the admissible-parameter conditions and the required smallness bounds for $(p_{\mathrm{match}},p_{\mathrm{spur}},p_{\mathrm{mean}},\eta_{0})$. In this section we give an explicit dense $C_{4}$-free PMRS scaffold and verify all of these quantities. A key technical point is the mean-control term $p_{\mathrm{mean}}$, which we bound by combining the Lipschitz-and-concentration Section 4.7 with an additional centering estimate (Section 4.7) that is available in this dense regime. This yields the $\Omega(\sqrt{n})$ adaptive one-sided lower bound.
$C_{4}$-free PMRS scaffolds from difference-free shifts.

### v2

@@ 1.3.3 Upper bounds
Bottleneck routing from a violation matching.
Our algorithmic upper bounds use the following combinatorial fact: distance to monotonicity is witnessed by a large matching of violated comparable pairs in the transitive closure. Concretely, if $f$ is $\varepsilon$-far, then the violation graph (on comparable pairs) contains a matching $M$ of size $|M|\geq\varepsilon n$ (via Konig’s theorem), so there are many disjoint source–sink violations $(u\!\rightsquigarrow\!v,\ f(u)>f(v))$. The difficulty is that these violations typically live in the transitive closure, denoted by $\mathrm{TC}(G)$, while the input provides only the transitive reduction (Hasse diagram), whose edge set can be much smaller and may contain very few direct violations. The key observation is that if direct violating edges are scarce, then they act as bottlenecks through which many matched violations are routed, forcing a large supply of violated pairs that can be detected by simple sampling.
The main combinatorial step is to partition the violation matching $M$ according to a first violating edge in the transitive reduction. Fix, for each $(u,v)\in M$, an arbitrary directed path $P_{uv}$ in the transitive reduction from $u$ to $v$. Since $f(u)>f(v)$, along that path there must be a first edge $(x,y)$ where the inequality flips, i.e., a violating edge of the transitive reduction. Assign $(u,v)$ to this first violating edge, and thereby partition
@@
This assignment has two consequences.
@@
    Because $M$ is a matching, the sets of left endpoints $\{A_{e}\}$ and right endpoints $\{B_{e}\}$ of the parts $M_{e}$ are disjoint across different bottlenecks $e$. This disjointness is the main technical reason the probabilistic analysis can be carried out globally.
From here, Section 5 develops two testers (Algorithms 1 and 2), each using this dichotomy in a different way and proved in Theorems 5.1 and 5.2.
@@
The first tester (Algorithm 1, see also Section 5.1) samples a moderate number of edges from the transitive reduction $E$ and, independently, a moderate number of comparable pairs from $E(\mathrm{TC}(G))$. The analysis uses the following dichotomy:
  • If many edges of the transitive reduction are violated (i.e., $|F_{f}|$ is large), then uniform sampling from $E$ finds a witness quickly.
  • If few edges of the transitive reduction are violated, then the bottleneck partition above implies a quadratic blow-up in the number of violated comparable pairs in the closure: each bottleneck edge $e$ with a large bucket $M_{e}$ certifies many violated pairs of the form $(a,b)$ with $a\in A_{e}$, $b\in B_{e}$. The median trick used in the proof formalizes this amplification: by splitting endpoints around a median threshold $t_{e}$, one obtains large subsets $A^{+}_{e}\subseteq A_{e}$, $B^{-}_{e}\subseteq B_{e}$ such that every $(a,b)\in A^{+}_{e}\times B^{-}_{e}$ is a violated pair in $\mathrm{TC}(G)$. Thus each $e$ contributes $\Omega(|M_{e}|^{2})$ violated comparable pairs, and summing (plus Cauchy–Schwarz) yields a lower bound of order $|M|^{2}/|F_{f}|$ on the number of violated closure edges.
@@
samples in each loop (up to constants). This is a tradeoff between easy local witnesses and many global witnesses: if one cannot see violations locally, they must manifest globally in abundance, and vice versa. This bound can beat the classical $O(\sqrt{n/\varepsilon})$ tester on sparse reachability structures, precisely when $m\ell=o(n^{3})$ for constant $\varepsilon$.
@@
The second tester (Algorithm 2, see also Section 5.2) removes dependence on $\ell$ altogether by avoiding uniform sampling from $\mathrm{TC}(G)$. Instead, after a first stage that samples edges of the transitive reduction (to catch the many-violating-edges case), it runs a second stage that samples two random multisets of vertices $Q_{L},Q_{R}$ and then checks all comparable pairs between them (reachability tests and scanning $Q_{L}\times Q_{R}$ are free in the model). Here the bottleneck partition yields a stronger conclusion: the disjoint endpoint structure means that if there are few violating edges, then the matching $M$ must be split across few buckets, so some bucket $M_{e}$ must be large; and if $Q_{L}$ hits $A^{+}_{e}$ and $Q_{R}$ hits $B^{-}_{e}$ for the same $e$, the tester certifiably rejects.
The technical obstacle in making this quantitative is dependence: sampling vertices with replacement creates complicated correlations when analyzing the event “for some $e$, we hit both sides of $A^{+}_{e}\times B^{-}_{e}$.” Section 5 resolves this with a standard Poissonization argument. By coupling the multinomial sampling counts with independent Poisson variables, the analysis turns the success events for different bottlenecks $e$ into (essentially) independent events, exactly because the $A^{+}_{e}$’s are disjoint and the $B^{-}_{e}$’s are disjoint. This produces a product bound and again leverages Cauchy–Schwarz to relate $\sum_{e}|M_{e}|^{2}$ to $|M|^{2}/|F_{f}|$. Choosing the stage sizes to balance the regime in which many edges of the transitive reduction are violated with the few-bottlenecks regime yields
@@ 1.4 Related work
Monotonicity testing on the Boolean hypercube was initiated by Goldreich et al. [21]. For product domains, subsequent work gave improved testers, near-optimal bounds for hypercubes and hypergrids, and lower bounds over product domains and hypergrids [11, 12, 16, 5, 10, 13, 14, 15]; connections to isoperimetric inequalities for the Boolean cube were developed in [26]. For general posets, testing was initiated by Fischer et al. [19], and a complementary line used transitive-closure spanners (TC-spanners) and reachability sparsification to design testers on structured graph families [7, 28]. Beyond standard testing, distance estimation and tolerant testing were studied in [1, 27], while local reconstruction of monotone functions and its limitations appear in [6, 31]. More recently, for real-valued functions on product domains, monotonicity has been connected to isoperimetric inequalities for ordered ranges [9, 12]. In contrast, for explicit-DAG monotonicity testing with real-valued ranges, prior lower bounds in both the non-adaptive and adaptive models remained far from the $\sqrt{n}$ scale.
@@
Graphs whose edges can be partitioned into induced matchings (Ruzsa–Szemerédi graphs) have been extensively studied in extremal graph theory since [30]. In the linear regime $r=cn$, the number of induced matchings $t$ is tightly constrained in several ranges of $c$: Fox, Huang, and Sudakov show that $t=O(1)$ for $c>1/4$, that $t=\Theta(\log n)$ for $c=1/4$, and that $t=O(n/\log n)$ for every fixed $1/5+\varepsilon\leq c<1/4$ [20]. On the other hand, for every fixed $c<1/4$, Fischer et al. construct graphs with induced matchings of size $cn$ and with $t=n^{\Omega(1/\log\log n)}$ such matchings, which underlie their Boolean monotonicity-testing lower bound on general posets [19]. If one allows slightly sublinear matchings instead, Alon, Moitra, and Sudakov construct nearly complete graphs whose edges decompose into induced matchings of size $n^{1-o(1)}$, with total edge count $\binom{n}{2}-o(n^{2})$ [3].
Decompositions into positive matchings have also been studied under the name of positive matching decompositions [18]. In that line of work, the natural question is whether the edge set of a graph can be partitioned into positive matchings, and if so, into how many parts. Concretely, positive matchings are characterized by the absence of $M$-alternating closed walks (where $M$ denotes the matching under consideration) on the $M$-saturated vertices [18]; in bipartite graphs this is equivalent to being alternating-cycle-free (uniquely restricted matching) [24]. For our purposes, PMRS is the real-valued analogue of RS packings on bipartite graphs, and positivity is exactly the condition that allows a matching to appear as the set of violated edges.
@@
Section 2 introduces notation and basic facts. Section 3 develops the non-adaptive two-sided lower bounds and explicit PMRS constructions. Section 4 proves adaptive one-sided lower bounds. Section 5 collects the algorithmic upper bounds parameterized by $m$ and $\ell$.
@@ 3.3 Non-adaptive two-sided lower bounds from PMRS
We begin by proving a non-adaptive two-sided lower bound from PMRS: for any $(s,4\varepsilon)$-PMRS graph, any non-adaptive two-sided tester requires $\Omega(\sqrt{s})$ queries. The proof builds a YES/NO distribution using a margin-normalized witness and a correlated-noise construction, which we set up below.
@@ 3.5 Main lower bound
We now combine the PMRS lower-bound argument with the explicit construction. As a first step, we state a refinement lemma that converts larger positive matchings into many smaller ones while preserving positivity.
@@ 3.6 Bounded ordered ranges and polylogarithmic ranges
The PMRS lower-bound argument does not inherently require unboundedness (two-sided).
The non-adaptive two-sided lower bound derived from an $(s,\varepsilon_{0})$-PMRS decomposition (Theorems 3.8 and 3.5) is information-theoretic: it constructs YES/NO distributions by adding a constant-alphabet noise to a slack monotone base and argues that a non-adaptive query set cannot distinguish the two distributions unless it queries both endpoints of some matching edge. The proof uses only the relative order of function values and a constant slack/margin separation; it does not rely on the range being unbounded beyond the ability to realize the prescribed pattern of inequalities. Consequently, the same argument extends to $[B]$ provided that the construction can be implemented with values lying in $[B]$ (after an order-preserving relabeling and an affine scaling/translation if needed).
@@ 4 Adaptive One-Sided Lower Bounds
The rest of this section is organized as follows. We first set up the PMRS base graph and truncated Gibbs distribution (Section 4.1), prove edge-tail bounds and parameter admissibility (Section 4.2), establish farness and control of spurious queried violations (Section 4.3), and bound adaptive information leakage via a Green-function comparison (Section 4.4). Next, we combine these ingredients with decoding and Fano’s inequality (Section 4.5) to obtain a fixed-budget lower bound (Section 4.6). Finally, we instantiate parameters on an explicit dense base graph to derive the $\Omega(\sqrt{n})$ adaptive one-sided lower bound (Section 4.7).
@@
(A) Global parameters and base-graph quantities.
@@
$n,n_{0}$ opening of this section total number of vertices and side size in the bipartite base graph ($n=2n_{0}$) throughout this section
@@
$\delta,\Delta,m$ opening definitions in Section 4.1 min/max degrees and edge count of the base graph Sections 4.2 and 4.7
$Q$ opening definitions in Section 4.1 query-budget scale for leakage and fixed-budget lower bounds Sections 4.4 and 4.6
@@
4.1 PMRS base graph and truncated Gibbs distribution
In this subsection, we define the truncated Gibbs distribution on a PMRS base graph and record the basic one-site curvature setup used later.
@@ 4.1 PMRS base graph and truncated Gibbs distribution
Step 5: bound the prefactor explicitly. Using $m_{v}=\alpha+2\lambda\deg(v)\geq 2\lambda\deg(v)$ and, since $\deg(v)\geq\delta$ and we assume $\alpha\leq(\lambda+\beta)\delta$, we have $M_{v}\leq\alpha+2(\lambda+\beta)\deg(v)\leq 3(\lambda+\beta)\deg(v)$, together with $A\geq c_{1}\deg(v)(\beta-\lambda)\gamma$, we obtain
@@ 4.2 Edge-tail bounds and admissible parameter set
Additional graph-theoretic assumption.
The next lemma (and only this lemma) uses an additional graph-theoretic property of the base graph: we assume that the underlying undirected graph of $U$ is $C_{4}$-free (has no $4$-cycles). This hypothesis is satisfied by the explicit family constructed in Section 4.7 below. All subsequent uses of $p_{\mathrm{drift}}(u)$ and the edge-tail bounds depend on this hypothesis.
@@
4.3 Farness and control of spurious queried violations
@@ 4.6 Budgeted one-sided lower bound
We now combine the bounds controlling spurious queried violations and leakage from the previous subsections to obtain a fixed-budget lower bound for adaptive one-sided testers.
Theorem 4.22 (Fixed-budget information lower bound).
@@
4.7 An $\Omega(\sqrt{n})$ instantiation and the mean-control bound
To apply the fixed-budget lower bound Theorem 4.22, we verify the admissible-parameter conditions and the required smallness bounds for $(p_{\mathrm{match}},p_{\mathrm{spur}},p_{\mathrm{mean}},\eta_{0})$. In this section we give an explicit dense $C_{4}$-free PMRS base graph and verify all of these quantities. To bound the mean-control term $p_{\mathrm{mean}}$, we combine the Lipschitz and concentration bounds in Section 4.7 with the additional centering estimate (Section 4.7) available in this dense regime. This yields the $\Omega(\sqrt{n})$ adaptive one-sided lower bound.
$C_{4}$-free PMRS base graphs from difference-free shifts.